OpenStudy (anonymous):

A baseball player hits a baseball into the outfield. The following equation models the ball's path. If not caught, when will the baseball land? Round to the nearest tenth second. h = -16t2 + 50t + 4 A. 3.0 seconds B. 3.2 seconds C. 3.4 seconds D. 3.6 seconds

OpenStudy (anonymous):

−16t2+50t+4=0 solve for t

OpenStudy (anonymous):

might want to start by dividing by -2

OpenStudy (anonymous):

8t^2−25t−2=0 use the quadratic formula

OpenStudy (anonymous):

8x^2-25x-2=0

OpenStudy (anonymous):

So the answer is A

OpenStudy (anonymous):

I mean B, sorry